On January 1, 2016, CIATS Corp. issued $100,000 face value, 9% bonds for $107,106 cash. The bonds are dated January 1, 2016, and mature on December 31, 2020. Interest will be paid semiannually on January 1 and July 1. The effective rate is 8% and CIATS uses the effective method to record the cash payments. What is the accrued interest amount that CIATS Corp. report in its December 312016 balance sheet?
Answer #1

Interest paid = 100000*9%*6/12 = 4500

Interest expense on July 1 = 107106*8%*6/12 = 4284

Premium amortization = 4284-4500 = 216

Interest expense on Dec 31 = (107106-216)*8%*6/12 = 4275.60 or 4276

So accrued interest amount = $4275.60 or $4276